OpenStudy (anonymous):

x-14 over x-1 = 4- 2x over x-1 =?

OpenStudy (anonymous):

\[x = 6\]

OpenStudy (anonymous):

\[(x-14)\div(x-1) = (4-2x)\div(x-1)\] same numerators go! (x = 1 cannot be an answer) \[x-14 = 4 - 2x \rightarrow 3x = 18 \rightarrow x = 6\]